OpenStudy (anonymous):

Can someone please help me? Kristin uses her credit card to purchase a new television for $863.35. She can pay off up to $425 per month. The card has an annual rate of 29.9% compounded monthly. How much will she pay in interest?

OpenStudy (anonymous):

To get your interest, just multiply $863.35 with $425, then multiply your product with 29.9% as a decimal.

OpenStudy (anonymous):

Why would you multiply 863.35 by 425?

OpenStudy (anonymous):

Well it's the formula for interest.

OpenStudy (anonymous):

oh.. So even though you are paying off 425 you still multiply it?

OpenStudy (anonymous):

Yup. Cause once you multiple 29.9% as a decimal it'll become less since the decimal point will be somewhere in the answer.

OpenStudy (anonymous):

That does not make sense. If you do that then 863.35*425=366923.75 than multiplied by .299 you get 109710.20 The answer choices are $41.28 $11.52 $34.12 $47.47

OpenStudy (anonymous):

Really? Let me double check and google the formula for interest. I'm sorry, but this is how I was taught in 8th grade.

OpenStudy (anonymous):

Yeah. I have never learned this before. so I am just attempting to figure it out.

OpenStudy (anonymous):

How many months are there?

OpenStudy (anonymous):

I do not know. I do not think that is the correct formula.

OpenStudy (anonymous):

Hmm.. Well google up interest and find your formula because that is the original interest formula. But I'm sorry if I can't help much, I;m a freshman. lol.

OpenStudy (anonymous):

I'm

OpenStudy (anonymous):

You can always contact your teacher although.

OpenStudy (anonymous):

so it would be 90.4%?

OpenStudy (anonymous):

my bad:\[1+interestRate_{month}=\sqrt[12]{1+interestRate_{year}}\]

OpenStudy (anonymous):

so 2.2%

OpenStudy (anonymous):

absolutely

OpenStudy (anonymous):

so you would do 863.35*.022=19.04 and than 863.35+19.04=882.39 than -425 would be 457.39

OpenStudy (anonymous):

Then I am lost. Or maybe I was already lost.. and I am doing something else worng.

OpenStudy (anonymous):

now you have to apply this interest rate to your left over balance at the end of each month, and then subtract whatever payment you can make

OpenStudy (anonymous):

exactally, do that until you've paid it off

OpenStudy (anonymous):

and than I would add all the interest together?

OpenStudy (anonymous):

i like the way you do it cause your keeping track of how much interest each month. yup

OpenStudy (anonymous):

10+19.04+.93= 29.97?

OpenStudy (anonymous):

But that is not an answer either.. So what did I do wrong?

OpenStudy (anonymous):

hmmm, i got 30.04 which is close enough to yours.give me a sec to go through it

OpenStudy (anonymous):

whom ever provided the answers found the monthly interest rate by dividing the yearly rate by 12. This yielded a monthly rate of 2.49%. Confusing, but how banks actually calculate their yearly interest rates to make them appear lower then they actually are (29.9% vs 34.35%) If you divide 29.9 by 12 instead of taking the 12th root of 1.299 and follow all your steps you'll get an appropriate answer

OpenStudy (anonymous):

21.50 = 459.86 11.45 = 46.31 1.15 so the answer should be 34.1

OpenStudy (anonymous):

So that is close enough to 34.12 so that is the answer.

OpenStudy (anonymous):

Thank you so much! You are amazing!

OpenStudy (anonymous):

i think the 12th root is more correct usually, maybe check with your teacher to see if they can be more specific in the question and get rid of the ambiguity

OpenStudy (anonymous):

Yeah. I am going to an online school. So all of the lessons are prewritten and since I am so far ahead in school there are times when the teachers have not even looked at the lessons I am on. But thank you for your help you have been amazing!